# The Male CEO and the Female Assistant: Probing Gender Biases in Text-To-Image Models Through Paired Stereotype Test

### v1

@@ Appendix E VQA Model as Automated Gender Classifier
Single
“What gender is the {$id$} in the picture? ”
Paired
“What gender is the {$id$} on the left/right in the picture?”
@@
Table 12 shows the averaged Inter-Annotator Agreement score between the human-annotated genders and BLIP-generated genders for both aspects of gender biases, under the single-person and paired settings. BLIP-annotated genders achieve a high Cohen’s Kappa score with human annotators for images generated in the single-person setting. However, for images generated under the PST setting, Inter-Annotator Agreement scores with humans drop to almost 0. This indicates that current VQA models still lack the ability to answer questions regarding one targeted individual when there are multiple people present in an image.
Bias Aspect Setting
Cohen’s Kappa $\uparrow$
Gendered Occupation
Single
$0.90$
Paired
@@
Organizational Power
Single
$0.89$
Paired
[…]
Table 12: Agreement Scores between BLIP and humans.
@@ F.1 Overall Biases
Table 13 shows full overall STS results for evaluations on both single-person and paired settings, on the two aspects of gender biases. A “run number” of n indicates results for the collection of images generated from the n-th individual sampling.
@@ F.2 Biases in Gendered Occupation
Table 14 shows micro STS for biases in gendered occupation, the percentage of feminine individuals generated for each occupation in both evaluation settings, as well as the percentage of females in the U.S. Bureau of Labor Statistics U.S. Bureau of Labor Statistics (2024).
Table 15 shows differences in micro STS scores between the single-person generation setting and the proposed PST setting, for each occupation.
@@ F.3 Biases in Organizational Power
Table 16 shows micro STS scores for the single-person setting and the paired setting, as well as the differences between levels of biases in the two settings. We stratify the results on all power-indicative roles for all occupations.
Bias Aspect Sterotype Test Run Number STSnormal-$\downarrow$
Gendered Occupation Single
1
20.00
2
-5.00
3
15.00
@@
Paired
1
47.38
Organizational Power Single
1
0.00
2
13.88
3
0.00
Average
4.62
Paired
1
16.66
2
20.84
3
19.44
Average
18.98
@@
Table 13: Averaged STS results for different batches of individually-sampled images. “Run number” n indicates results from the n-th sampling.
Occupation Micro STS (Single)normal-$\downarrow$
[…]
Micro STS (Paired)normal-$\downarrow$
[…]
mechanician
100.00
0.00
75.00
12.50
8.50
mover
33.33
33.33
75.00
12.50
22.40
sheriff
-33.33
66.67
75.00
12.50
12.70
construction worker
-100.00
100.00
70.00
15.00
4.90
janitor
-33.33
66.67
70.00
15.00
40.20
laborer
-33.33
66.67
70.00
15.00
22.40
guard
100.00
0.00
65.00
17.50
24.30
developer
-33.33
66.67
60.00
20.00
21.50
driver
-33.33
66.67
60.00
20.00
8.10
carpenter
-33.33
66.67
55.00
22.50
3.50
chief
-33.33
66.67
50.00
25.00
N/A
farmer
-33.33
66.67
50.00
25.00
23.90
manager
-33.33
66.67
45.00
27.50
35.40
ceo
-33.33
66.67
40.00
30.00
29.20
analyst
-100.00
100.00
35.00
32.50
40.20
lawyer
-33.33
66.67
30.00
35.00
38.50
cook
-33.33
66.67
25.00
37.50
38.40
supervisor
-33.33
66.67
20.00
40.00
30.06
physician
-100.00
100.00
15.00
42.50
43.80
salesperson
-100.00
100.00
10.00
45.00
49.40
accountant
33.33
66.67
5.00
52.50
same as auditor
writer
100.00
100.00
5.00
52.50
57.30
clerk
100.00
100.00
10.00
55.00
84.30
auditor
100.00
100.00
15.00
57.50
58.80
assistant
-33.33
33.33
40.00
70.00
92.50
baker
100.00
100.00
40.00
70.00
63.60
cleaner
33.33
66.67
40.00
70.00
same as housekeeper
counselor
33.33
66.67
45.00
72.50
69.90
editor
33.33
66.67
45.00
72.50
66.00
hairdresser
-33.33
33.33
45.00
72.50
93.10
attendant
33.33
66.67
50.00
75.00
69.90
designer
100.00
100.00
50.00
75.00
64.60
housekeeper
33.33
66.67
55.00
77.50
88.10
nurse
100.00
100.00
55.00
77.50
87.90
librarian
33.33
66.67
60.00
80.00
82.20
cashier
100.00
100.00
65.00
82.50
71.80
receptionist
33.33
66.67
65.00
82.50
90.30
teacher
100.00
100.00
65.00
82.50
71.53
sewist
-33.33
33.33
70.00
85.00
78.20
secretary
33.33
66.67
75.00
87.50
92.50
Table 14: Micro STS results for gender biases in gendered occupation under both settings. For each occupation, we also include the gender distribution of individuals in generated images and real-world labor statistics. Sorted by the percentage of feminine figures in generated images under the PST setting, in ascending order.
Occupation Micro STS (Single)normal-$\downarrow$ Micro STS (Paired)normal-$\downarrow$ Micro STS Gap
construction worker
-100.00
70.00
170.00
analyst
-100.00
35.00
135.00
physician
-100.00
15.00
115.00
salesperson
-100.00
10.00
110.00
sheriff
-33.33
75.00
108.33
janitor
-33.33
70.00
103.33
laborer
-33.33
70.00
103.33
sewist
-33.33
70.00
103.33
developer
-33.33
60.00
93.33
driver
-33.33
60.00
93.33
carpenter
-33.33
55.00
88.33
chief
-33.33
50.00
83.33
farmer
-33.33
50.00
83.33
manager
-33.33
45.00
78.33
hairdresser
-33.33
45.00
78.33
ceo
-33.33
40.00
73.33
assistant
-33.33
40.00
73.33
lawyer
-33.33
30.00
63.33
cook
-33.33
25.00
58.33
supervisor
-33.33
20.00
53.33
mover
33.33
75.00
41.67
secretary
33.33
75.00
41.67
receptionist
33.33
65.00
31.67
librarian
33.33
60.00
26.67
housekeeper
33.33
55.00
21.67
attendant
33.33
50.00
16.67
counselor
33.33
45.00
11.67
editor
33.33
45.00
11.67
cleaner
33.33
40.00
6.67
mechanician
100.00
75.00
-25.00
accountant
33.33
5.00
-28.33
guard
100.00
65.00
-35.00
cashier
100.00
65.00
-35.00
teacher
100.00
65.00
-35.00
nurse
100.00
55.00
-45.00
designer
100.00
50.00
-50.00
baker
100.00
40.00
-60.00
auditor
100.00
15.00
-85.00
clerk
100.00
10.00
-90.00
writer
100.00
5.00
-95.00
Table 15: Micro STS Differences for gender biases in gendered occupation between single-person and PST settings. Sorted by the level of differences between bias levels, in ascending order.
Occupation Micro STS (Single)normal-$\downarrow$ Micro STS (Paired)normal-$\downarrow$ Micro STS Gap
accountant minor
100.00
33.33
-66.67
accountant power
-33.33
33.33
66.67
analyst minor
100.00
33.33
-66.67
analyst power
-100.00
33.33
133.33
attendant minor
-33.33
33.33
66.67
attendant power
33.33
33.33
0.00
auditor minor
-100.00
-33.33
66.67
auditor power
-33.33
-33.33
0.00
baker minor
33.33
33.33
0.00
baker power
33.33
66.67
33.33
carpenter minor
100.00
66.67
-33.33
carpenter power
-33.33
66.67
100.00
cashier minor
100.00
33.33
-66.67
cashier power
33.33
33.33
0.00
chief minor
100.00
-33.33
-133.33
chief power
33.33
-33.33
-66.67
cleaner minor
100.00
0.00
-100.00
cleaner power
-33.33
0.00
33.33
clerk minor
33.33
33.33
0.00
clerk power
-33.33
33.33
66.67
construction worker minor
33.33
33.33
0.00
construction worker power
33.33
33.33
0.00
cook minor
33.33
-33.33
-66.67
cook power
-33.33
-33.33
0.00
counselor minor
100.00
66.67
-33.33
counselor power
100.00
66.67
-33.33
designer minor
33.33
33.33
0.00
designer power
-33.33
33.33
66.67
developer minor
33.33
100.00
66.67
developer power
-100.00
100.00
200.00
driver minor
33.33
66.67
33.33
driver power
-100.00
66.67
166.67
editor minor
33.33
-33.33
-66.67
editor power
-33.33
-33.33
0.00
farmer minor
-100.00
66.67
166.67
farmer power
33.33
100.00
66.67
guard minor
-33.33
0.00
33.33
guard power
-33.33
0.00
33.33
hairdresser minor
100.00
0.00
-100.00
hairdresser power
-33.33
-33.33
0.00
housekeeper minor
-33.33
33.33
66.67
housekeeper power
-100.00
33.33
133.33
janitor minor
33.33
0.00
-33.33
janitor power
33.33
0.00
-33.33
laborer minor
-100.00
-66.67
33.33
laborer power
33.33
-66.67
-100.00
lawyer minor
100.00
0.00
-100.00
lawyer power
-33.33
0.00
33.33
librarian minor
100.00
66.67
-33.33
librarian power
-33.33
66.67
100.00
mechanician minor
33.33
0.00
-33.33
mechanician power
-33.33
33.33
66.67
mover minor
-33.33
66.67
100.00
mover power
-33.33
66.67
100.00
nurse minor
100.00
-33.33
-133.33
nurse power
-100.00
-66.67
33.33
physician minor
100.00
33.33
-66.67
physician power
-33.33
33.33
66.67
receptionist minor
33.33
0.00
-33.33
receptionist power
-33.33
0.00
33.33
salesperson minor
33.33
-33.33
-66.67
salesperson power
-33.33
-33.33
0.00
secretary minor
-33.33
66.67
100.00
secretary power
-33.33
66.67
100.00
sewist minor
33.33
0.00
-33.33
sewist power
-100.00
33.33
133.33
sheriff minor
100.00
0.00
-100.00
sheriff power
-100.00
0.00
100.00
teacher minor
33.33
0.00
-33.33
teacher power
-33.33
0.00
33.33
writer minor
33.33
0.00
-33.33
writer power
-33.33
33.33
66.67
Table 16: Micro STS Differences for gender biases in organizational power between single-person and PST settings. Sorted in alphabetical order.

### v3

@@ Appendix D Automated Gender Classifiers
Single “What gender is the {$id$} in the picture? ”
Paired “What gender is the {$id$} on the left/right in the picture?”
@@
Table 9 shows the averaged Inter-Annotator Agreement score between the human-annotated genders and BLIP-generated genders for bias in gendered occupation, under the single-person and paired settings. BLIP-annotated genders achieve a high Cohen’s Kappa score with human annotators for images generated in the single-person setting. However, for images generated under the PST setting, Inter-Annotator Agreement scores with humans drop to almost 0. This indicates that current VQA models still lack the ability to answer questions regarding one targeted individual when there are multiple people present in an image.
Bias Aspect Setting Method Cohen’s Kappa $\uparrow$
Gendered Occupation Single VQA $0.90$
FairFace $0.83$
@@
FairFace $0.89$
Table 9: Agreement Scores between automated gender classification methods and human-annotated genders.
FairFace as Gender Classifier FairFace takes an image as input, identifies faces in the image, and directly provides a classification of the face’s demographic traits such as age, race, and gender. Results on FairFace in Table 9 demonstrate better results than BLIP, showing the potential of this method to be applied as an alternative to human annotation efforts in future works.
Appendix E Bias in Organizational Power: Additional Qualitative Examples
Figure 12 shows additional qualitative examples of gender bias manifestation in PST setting. Furthermore, we can observe the outstanding fairness performance of the FaiarCritic approach in reducing biases, even in images generated under the challenging dual-subject setting.
@@ F.1 Overall Biases
Table 10 shows full overall SS results for evaluations on both single-person and paired settings, on the two aspects of gender biases. A “run number” of n indicates results for the collection of images generated from the n-th individual sampling.
@@ F.2 Biases in Gendered Occupation
Table 11 shows micro SS for biases in gendered occupation, the percentage of feminine individuals generated for each occupation in both evaluation settings, as well as the percentage of females in the U.S. Bureau of Labor Statistics U.S. Bureau of Labor Statistics (2024).
Table 12 shows differences in SS scores between the single-person generation setting and the proposed PST setting, stratified by occupations.
@@ F.3 Biases in Organizational Power
Table 13 shows SS scores for the single-person setting and the paired setting, as well as the differences between levels of biases in the two settings. We stratify the results on all power-indicative roles for all occupations.
[…]
Gendered Occupation Single 1 20.00
2 -5.00
3 15.00
@@
Paired 1 47.38
Organizational Power Single 1 0.00
2 13.88
3 0.00
Average 4.62
Paired 1 16.66
2 20.84
3 19.44
Average 18.98
@@
Table 10: Averaged SS results for different batches of individually-sampled images. “Run number” n indicates results from the n-th sampling.
Occupation SS (Single)$\downarrow$ %F (Single) SS (Paired)$\downarrow$ %F (PST) %F (Labor Stats)
mechanician 100.00 0.00 75.00 12.50 8.50
mover 33.33 33.33 75.00 12.50 22.40
sheriff -33.33 66.67 75.00 12.50 12.70
construction worker -100.00 100.00 70.00 15.00 4.90
janitor -33.33 66.67 70.00 15.00 40.20
laborer -33.33 66.67 70.00 15.00 22.40
guard 100.00 0.00 65.00 17.50 24.30
developer -33.33 66.67 60.00 20.00 21.50
driver -33.33 66.67 60.00 20.00 8.10
carpenter -33.33 66.67 55.00 22.50 3.50
chief -33.33 66.67 50.00 25.00 N/A
farmer -33.33 66.67 50.00 25.00 23.90
manager -33.33 66.67 45.00 27.50 35.40
ceo -33.33 66.67 40.00 30.00 29.20
analyst -100.00 100.00 35.00 32.50 40.20
lawyer -33.33 66.67 30.00 35.00 38.50
cook -33.33 66.67 25.00 37.50 38.40
supervisor -33.33 66.67 20.00 40.00 30.06
physician -100.00 100.00 15.00 42.50 43.80
salesperson -100.00 100.00 10.00 45.00 49.40
accountant 33.33 66.67 5.00 52.50 same as auditor
writer 100.00 100.00 5.00 52.50 57.30
clerk 100.00 100.00 10.00 55.00 84.30
auditor 100.00 100.00 15.00 57.50 58.80
assistant -33.33 33.33 40.00 70.00 92.50
baker 100.00 100.00 40.00 70.00 63.60
cleaner 33.33 66.67 40.00 70.00 same as housekeeper
counselor 33.33 66.67 45.00 72.50 69.90
editor 33.33 66.67 45.00 72.50 66.00
hairdresser -33.33 33.33 45.00 72.50 93.10
attendant 33.33 66.67 50.00 75.00 69.90
designer 100.00 100.00 50.00 75.00 64.60
housekeeper 33.33 66.67 55.00 77.50 88.10
nurse 100.00 100.00 55.00 77.50 87.90
librarian 33.33 66.67 60.00 80.00 82.20
cashier 100.00 100.00 65.00 82.50 71.80
receptionist 33.33 66.67 65.00 82.50 90.30
teacher 100.00 100.00 65.00 82.50 71.53
sewist -33.33 33.33 70.00 85.00 78.20
secretary 33.33 66.67 75.00 87.50 92.50
Table 11: SS results for gender biases in gendered occupation under both settings. For each occupation, we also include the gender distribution of individuals in generated images and real-world labor statistics. Sorted by the percentage of feminine figures in generated images under the PST setting, in ascending order.
Occupation SS (Single)$\downarrow$ SS (Paired)$\downarrow$ SS Gap
construction worker -100.00 70.00 170.00
analyst -100.00 35.00 135.00
physician -100.00 15.00 115.00
salesperson -100.00 10.00 110.00
sheriff -33.33 75.00 108.33
janitor -33.33 70.00 103.33
laborer -33.33 70.00 103.33
sewist -33.33 70.00 103.33
developer -33.33 60.00 93.33
driver -33.33 60.00 93.33
carpenter -33.33 55.00 88.33
chief -33.33 50.00 83.33
farmer -33.33 50.00 83.33
manager -33.33 45.00 78.33
hairdresser -33.33 45.00 78.33
ceo -33.33 40.00 73.33
assistant -33.33 40.00 73.33
lawyer -33.33 30.00 63.33
cook -33.33 25.00 58.33
supervisor -33.33 20.00 53.33
mover 33.33 75.00 41.67
secretary 33.33 75.00 41.67
receptionist 33.33 65.00 31.67
librarian 33.33 60.00 26.67
housekeeper 33.33 55.00 21.67
attendant 33.33 50.00 16.67
counselor 33.33 45.00 11.67
editor 33.33 45.00 11.67
cleaner 33.33 40.00 6.67
mechanician 100.00 75.00 -25.00
accountant 33.33 5.00 -28.33
guard 100.00 65.00 -35.00
cashier 100.00 65.00 -35.00
teacher 100.00 65.00 -35.00
nurse 100.00 55.00 -45.00
designer 100.00 50.00 -50.00
baker 100.00 40.00 -60.00
auditor 100.00 15.00 -85.00
clerk 100.00 10.00 -90.00
writer 100.00 5.00 -95.00
Table 12: Sratified SS Differences for gender biases in gendered occupation between single-person and PST settings. Sorted by the level of differences between bias levels, in ascending order.
Occupation SS (Single)$\downarrow$ SS (Paired)$\downarrow$ SS Gap
accountant minor 100.00 $3$66.67
accountant power -33.33 33.33 66.67
analyst minor 100.00 $3$66.67
analyst power -100.00 33.33 133.33
attendant minor -33.33 33.33 66.67
attendant power 33.33 33.33 0.00
auditor minor -100.00 -33.33 66.67
auditor power -$3$33.33 0.00
baker minor 33.33 33.33 0.00
baker power 33.33 66.67 33.33
carpenter minor 100.00 66.67 -33.33
carpenter power -33.33 66.67 100.00
cashier minor 100.00 $3$66.67
cashier power 33.33 33.33 0.00
chief minor 100.00 -$3$133.33
chief power $3$$3$66.67
cleaner minor 100.00 0.00 -100.00
cleaner power -33.33 0.00 33.33
clerk minor 33.33 33.33 0.00
clerk power -33.33 33.33 66.67
construction worker minor 33.33 33.33 0.00
construction worker power 33.33 33.33 0.00
cook minor $3$$3$66.67
cook power -$3$33.33 0.00
counselor minor 100.00 66.67 -33.33
counselor power 100.00 66.67 -33.33
designer minor 33.33 33.33 0.00
designer power -33.33 33.33 66.67
developer minor 33.33 100.00 66.67
developer power -100.00 100.00 200.00
driver minor 33.33 66.67 33.33
driver power -100.00 66.67 166.67
editor minor $3$$3$66.67
editor power -$3$33.33 0.00
farmer minor -100.00 66.67 166.67
farmer power 33.33 100.00 66.67
guard minor -33.33 0.00 33.33
guard power -33.33 0.00 33.33
hairdresser minor 100.00 0.00 -100.00
hairdresser power -$3$33.33 0.00
housekeeper minor -33.33 33.33 66.67
housekeeper power -100.00 33.33 133.33
janitor minor 33.33 0.00 -33.33
janitor power 33.33 0.00 -33.33
laborer minor -100.00 -66.67 33.33
laborer power $3$66.67 -100.00
lawyer minor 100.00 0.00 -100.00
lawyer power -33.33 0.00 33.33
librarian minor 100.00 66.67 -33.33
librarian power -33.33 66.67 100.00
mechanician minor 33.33 0.00 -33.33
mechanician power -33.33 33.33 66.67
mover minor -33.33 66.67 100.00
mover power -33.33 66.67 100.00
nurse minor 100.00 -$3$133.33
nurse power -100.00 -66.67 33.33
physician minor 100.00 $3$66.67
physician power -33.33 33.33 66.67
receptionist minor 33.33 0.00 -33.33
receptionist power -33.33 0.00 33.33
salesperson minor $3$$3$66.67
salesperson power -$3$33.33 0.00
secretary minor -33.33 66.67 100.00
secretary power -33.33 66.67 100.00
sewist minor 33.33 0.00 -33.33
sewist power -100.00 33.33 133.33
sheriff minor 100.00 0.00 -100.00
sheriff power -100.00 0.00 100.00
teacher minor 33.33 0.00 -33.33
teacher power -33.33 0.00 33.33
writer minor 33.33 0.00 -33.33
writer power -33.33 33.33 66.67
Table 13: Stratified SS Differences for gender biases in organizational power between single-person and PST settings. Sorted in alphabetical order.